OpenStudy (anonymous):

I will post the question using the equation editor. x 1/5 x 2/10 x 1/25 x 2/5

OpenStudy (anonymous):

\[(x \left(\begin{matrix}1 \\ 5\end{matrix}\right) )(x \left(\begin{matrix}1 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

its 1/5